OpenStudy (anonymous):

2^(n+201) + 2^(n+201)= 2^(n+202) is this possible or just a mistake?

OpenStudy (angela210793):

=2*2^(n+201)

OpenStudy (dumbcow):

no that is correct 2*2^(n+201) = 2^(n+201)+1 = 2^(n+202)

OpenStudy (anonymous):

Yes it is. \[2^{n +201} + 2^{201 +n} = 2(2^{201+n})\]\[x^n .x^b = x^{b+n}\]\[2^{201+1+n} = 2^{202 +n}\]

OpenStudy (anonymous):

i see it and am still not sure...where is the 1 coming from? do you always add a 1 if there are two of the same numbers?

OpenStudy (dumbcow):

The one came from the 2 on the outside..it has an exponent of 1 right? when you multiply the same base, you add the exponents \[\rightarrow 2^{1}*2^{n+201} = 2^{n+201+1} = 2^{n+202}\]

OpenStudy (anonymous):

perfecto:D Thanx!

OpenStudy (dumbcow):

no problem